OpenStudy (anonymous):

PLEASE HELP :3 Solve 2x + 5y = −13 3x − 4y = −8 (4, 1) (−4, 1) (4, −1) (−4, −1)

OpenStudy (anonymous):

@Jack1 for ones like these: there is no parenthesis, or exponents, so I staret with the multiplication.. I want to get my y's to one side.. how do I do that? I divide them?

OpenStudy (jack1):

i like to subtract first to isolate the one term on one side so 2x + 5y = −13 2x = -13 -5y x = -(13/2) - (5/2)y

OpenStudy (jack1):

or if not, get them both in terms of 6x

OpenStudy (jack1):

2x + 5y = -13 so 6x + 15y = -39 and 3x - 4y = -8 so 6x - 8y = -16 therefore 6x = -39 - 15y and 6x = -16 + 8y

OpenStudy (anonymous):

AH I thought so, u do them apart and I knew I took the 5 from that side

OpenStudy (nurali):

2x + 5y = −13-------(1) 3x − 4y = −8 --------(2) in equation 1 2x=-13-5y x=-13-5y/2 put x in equation 2 3(-13-5y/2)-4y=-8 -39-15y/2-4y=-8 -39-15y-8y=-16 -23y=-16+39 -23y=23 y=-1 put y in equation 1 2x + 5y = −13 2x+5(-1)=-13 2x-5=-13 2x=-13+5 2x=-8 x=-4

OpenStudy (jack1):

6x = -39 - 15y = -16 + 8y -39 - 15y = -16 + 8y -39 = -16 + 23y -23 = 23y -1 = y

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

The next one I have I have to get the value of y.. is it the same way we get the value of x?

OpenStudy (jack1):

exactly the same way, it doesn't matter if the term is called x or y or q, its all the same when using simultaneous equations

OpenStudy (anonymous):

Ok I think I gotcha

OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x − 3y = −11 2x − 6y = −14 −1 0 1 2

OpenStudy (anonymous):

This is what I have. ill let u no what I get

OpenStudy (jack1):

cools

OpenStudy (anonymous):

ok so I got

OpenStudy (anonymous):

5x-3y=-11 +3y ____________ 5x=-11+3y

OpenStudy (anonymous):

correct?

OpenStudy (uri):

where is +3y coming from?

OpenStudy (jack1):

yep

OpenStudy (jack1):

ignore uri, he's a little slow and apparently i hate him ;D

OpenStudy (anonymous):

Oh..

OpenStudy (uri):

Huh? I'm a girl..and that's not the way to talk.

OpenStudy (anonymous):

Where do I go from 5x=-11+3y

OpenStudy (jack1):

sure u are... online anyway an i bet u tell that to all the boys ;P

OpenStudy (uri):

Ah? last time i checked..i was a girl.

OpenStudy (jack1):

double it change it to 10x

OpenStudy (anonymous):

..... sooo no one is going to help me but we will argue on my question?... Thanks guys

OpenStudy (anonymous):

Now your confusing me and not helping me

OpenStudy (jack1):

5x=-11+3y 10x = -22 + 6

OpenStudy (uri):

You are not even teaching her the right thing.Also,You are very mean and you don't know how to respond to someone..when they ask you a question. I might link you this: http://openstudy.com/code-of-conduct Next time if you're mean to me.I'll report you to Preetha. @Jack1

OpenStudy (jack1):

so thats equation 1 in a form we can use now for equation 2: 2x − 6y = −14 change it so u get 10x on one side

OpenStudy (jack1):

sorry, back to equation 1: it should be 10x = -22 + 6y, not 10x = -22 + 6, my bad now equation 2: 2x − 6y = −14 (times 5) 10x - 30y = -70 (add 30y to both sides) 10x - 30y + 30y = -70 + 30y 10x = -70 + 30y

OpenStudy (jack1):

so equation 1 is now 10x = -22 + 6y and equation 2 is now 10x = -70 + 30y so 10x = -22 + 6y = -70 + 30y -22 + 6y = -70 + 30y can u solve from here?

thomaster (thomaster):

@Jack1 She made a new question http://openstudy.com/study#/updates/52120b0fe4b0450ed75d4d22

OpenStudy (jack1):

@raebaby420

OpenStudy (uri):

Huehue.

OpenStudy (jack1):

god u are such a girl XP

OpenStudy (uri):

You called me a boy dat mad me cry..:3

OpenStudy (uri):

made*

OpenStudy (jack1):

im sorry i called u a boy, i was jus playing and i took it too far, my bad

OpenStudy (uri):

It's alright.